Please help me set up the equation for this word problem so that I can then solve.
Working together, Paul and Daniel can pick forty bushels of apples in 4.95 hours. Had he done it alone it would have taken Daniel 9 hours. Find how long it have taken Daniel 9 hours. Find how long it would take Paul to do it alone.

Answer by checkley79(3341)

1/9+1/P=1/4.95
(P+9)/9P=1/4.95 CROSS MULTIPLY.
9P=4.95(P+9)
9P=4.95P+44.55
9P-4.95P=44.95
4.05P=44.95
P=44.95/4.05
P=11.1 HOURS FOR PAUL TO PICK 40 BUSHELS OF APPLES ALONE.
PROOF:
1/9+1/11.1=1/4.95
(11.1+9)/9*11.1=1/4.95
20.1/99.9=1/4.95
99.9=20.1*4.95
99.9~99.5